OpenStudy (anonymous):

a dirt pile is cone shaped with a height of 10 feet and a diameter of 24 feet what is the volume?

OpenStudy (arindameducationusc):

use voume of cone formula

OpenStudy (arindameducationusc):

1/3pi*r^2*h where r=diameter/2=24/2

OpenStudy (arindameducationusc):

h=10

OpenStudy (zzr0ck3r):

\(V=\dfrac{1}{3}\pi r^2h\)

HanAkoSolo (jamierox4ev3r):

\(\huge V=\pi r^{2}\frac{h}{3}\) To make it more clear :) h=10 if your diameter is 24, divide that in half to get the radius r=12

HanAkoSolo (jamierox4ev3r):

\(\huge V=\dfrac{1}{3}\pi r^2h\) ^that can be used as well. Now plug in the values that you do know (h and r) and solve for V (volume).